OpenStudy (turingtest):

Fun integral that seems to have gotten deleted, so I'm posting it again.\[\int_0^\pi\frac{x\sin x}{1+\cos^2x}dx\](yes I am an idiot, I accidentally closed the other post)

OpenStudy (kainui):

Haha do I do trig substitution on this? I can't resist must go for it.

hero (hero):

My calc could do this in two seconds

OpenStudy (turingtest):

actually the answer on the test is given, ad I think wolf can do it to the problem on the test is to prove the result is what it is (it's from the entrance exam to a university here in Mexico)

OpenStudy (turingtest):

and*

OpenStudy (turingtest):

y'all wanna know the answer? the proof still remains...

hero (hero):

Torturing students at U of M I see

OpenStudy (turingtest):

Cinvestav, for masters and doctorates only

OpenStudy (anonymous):

\[I=\int_0^\pi\frac{x\sin x}{1+\cos^2x}\text{d}x\] \[x=\pi-t\]\[I=\int_\pi^0\frac{(\pi-t)\sin t}{1+\cos^2t}(-\text{d}t)=\int_0^\pi\frac{(\pi-t)\sin t}{1+\cos^2t}\text{d}t\]\[=\int_0^\pi\frac{\pi\sin t}{1+\cos^2t}\text{d}t-\int_0^\pi\frac{t\sin t}{1+\cos^2t}\text{d}t\]\[2I=\int_0^\pi\frac{\pi\sin t}{1+\cos^2t}\text{d}t=-\pi \ \tan^{-1} (\cos t) |_0^\pi=\frac{\pi^2}{4}\]\[I=\frac{\pi^2}{8}\]

OpenStudy (anonymous):

emm...\[2I=\int_0^\pi\frac{\pi\sin t}{1+\cos^2t}\text{d}t=-\pi \ \tan^{-1} (\cos t) |_0^\pi=\frac{\pi^2}{2}\]\[I=\frac{\pi^2}{4}\]

OpenStudy (turingtest):

@mukushla nailed it, as usual! Very similar to my approach, hats off :D

Parth (parthkohli):

How do you people do that? :|

Parth (parthkohli):

Stop being geniuses, or I'd kill myself one day for my ignorance. :P

OpenStudy (turingtest):

It's called practice, as I always say...

OpenStudy (kainui):

This just seems like guessing what the solutions to a quadratic equation are or what the integrating factor of an exact differential equation is without actually working it out. How do you practice this? I don't remember this in Cal 1 or 2, but I'm damn interested. Maybe I'm not conceptualizing integrals in their simplest possible form... if that makes sense.

OpenStudy (turingtest):

The best reply to that is 3 hints... hint one: integration by parts let me know how that goes and I will give more hints if necessary